| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  128  
 | 
08 Apr 2023, 10:29 | 
Andreas Galler | PRC | Issue | Changed EPS Interlock condition 
for SPB_XTD9_IMGPI/SWITCH/SCREEN_OUT
 
 
SPB/SFX called DOC this morning | 
   |  | 
  127  
 | 
07 Apr 2023, 23:57 | 
Yifeng Jiang | FXE | Issue | After receiving beam from SPB, we noticed 
that there is a werid shape of beam in the 
FEL imager which was not observed yesterday. | 
    |  | 
  126  
 | 
07 Apr 2023, 23:03 | 
Johan Bielecki | SPB | Shift summary | Beam delivery for proposal 3378
 
 
Achievements/Observations
 
 
 
	Collected data on liquid | 
   |  | 
  125  
 | 
07 Apr 2023, 23:00 | 
Alejandro Laso Garcia | HED | HED Shift Summary 07.04. | Morning - Users from proposal 3455 continued 
with their measurement program. Beam stable 
during the measurements. | 
   |  | 
  124  
 | 
04 Apr 2023, 11:18 | 
Johan Bielecki |   | XTD2 Si500 attenuator broken | Attachment 1: Beam in SPB hutch with Si500
 
 
Attachment 2: Beam without Si500 | 
     |  | 
  123  
 | 
04 Apr 2023, 08:37 | 
Frederik Wolff-Fabris | PRC | Issue | TUE 08:30AM:
 
 
The SA3 Cell 17 & 18 Phase 
shifters do not move and encoders are damaged. A | 
   |  | 
  122  
 | 
04 Apr 2023, 07:11 | 
Lisa Randolph, Carsten Baehtz, Alexander Pelka | HED | Summary day shift 2023-04-04 | 04 Apr 2023, 07:44 we realize that the 
SetpointsHED Macro lost all values, probably 
during the Karabo Crash. Since it is a macro, | 
    |  | 
  121  
 | 
03 Apr 2023, 06:38 | 
Loic, Yi-Ping | SCS | Night shift summary 02.04.2023 | Achievements:
 
 
Collected more scans on cyanocobalamin, 
both delays scans and dynamic XAS | 
   |  | 
  120  
 | 
02 Apr 2023, 04:46 | 
Benjamin van Kuiken | SCS | Late Shift Summary | Achievement
 
 
 
	Measured tXAS at the Fe 
L-edge of Fe(bpy)3 using both delay and energy | 
   |  | 
  119  
 | 
01 Apr 2023, 06:51 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	1-202 pulses, 564 kHz, 
stable delivery
 
	Beam was down | 
   |  | 
  118  
 | 
31 Mar 2023, 06:59 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	1-202 pulses, 564, 
282, 141 , 70.5 kHz, stable delivery
 
 
 
Optical | 
   |  | 
  117  
 | 
31 Mar 2023, 05:10 | 
Benjamin, Yi-Ping, Ru-Pan, Andreas | SCS | Late Shift Summary | Achievements
 
 
 
	Optimized BOZ for Co and 
Fe
 
	Estabilished Stable liquid | 
   |  | 
  116  
 | 
31 Mar 2023, 01:51 | 
Jörg Hallmann | MID | shift summary March 30th | - seeding at ~800 uJ
 
 
- repaired the laser interlock 
of A.08 | 
   |  | 
  115  
 | 
30 Mar 2023, 06:53 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	1-202 pulses, 564, 
282, 141 kHz, stable delivery
 
 
 
Optical | 
   |  | 
  114  
 | 
30 Mar 2023, 04:16 | 
Benjamin, Robert, Yi-Ping, Ru-Pan | SCS | Shift Summary | Achievements
 
 
 
	Found coarse timing between 
XFEL and optical laser
 
 
 
Problems | 
   |  | 
  113  
 | 
30 Mar 2023, 02:37 | 
Jörg Hallmann | MID | shift summary March 29th | achievments: alignment of the SDL (both 
branches)
 
 
measuring of samples & calibrants | 
   |  | 
  112  
 | 
29 Mar 2023, 01:31 | 
Jörg Hallmann | MID | Shift summary March 28th | Seeding at 9 keV with 800 - 1000uJ intensity
 
 
1. alignment of the lower branch 
of the SDL | 
   |  | 
  111  
 | 
28 Mar 2023, 15:26 | 
Angel Rodriguez Fernandez | MID | Issue | The door from MID side entrance was smashed 
and we lost the radiation permit.
 
 
The acciedent happened between | 
    |  | 
  110  
 | 
26 Mar 2023, 23:27 | 
Jayanath Koliyadu | SPB | Shift end | X-ray delivery
 
 
 
	1-165 pulses, 1.1 MHz - 
56.4 kHz, upto 350  uJ @ 50 pC
 
 
 
Optical | 
   |  | 
  109  
 | 
26 Mar 2023, 13:26 | 
Naresh Kujala | PRC | Status | Machine RC Raimond informed PRC that 
there is an issue with the A4 module that 
needs to fix to have a continuous operation. | 
   |   
 |